A car travelling at 30m/seconds brought to rest in a distance of 50 meters

1. calculate it acceleration 

2. time taken for it to come to rest 


1
Expert's answer
2021-03-25T19:28:50-0400

1)

"v^2=v_0^2+2as,""0=v_0^2+2as,""a=\dfrac{-v_0^2}{2s}=\dfrac{-(30\ \dfrac{m}{s})^2}{2\cdot50\ m}=-9\ \dfrac{m}{s^2}."

The sign minus means that the car decelerates.

2)

"v=v_0+at,""0=v_0+at,""t=\dfrac{-v_0}{a}=\dfrac{-30\ \dfrac{m}{s}}{-9\ \dfrac{m}{s^2}}=3.33\ s."